# Partonic Entropy of the Proton from DGLAP Evolution

### v2

@@ II Partonic entropy
Following the Gibbs-Shanon formula for entropy of a probability distribution $\{P_{n}\}$, defined on a discrete set of states nn [^1] In quantum mechanics, $S$ is the von Neumann entropy, $S=-{\rm Tr\,}\rho\ln\rho$, computed in the basis that diagonalizes the density operator $\rho$. ,
@@ V Relative entropy
which can be easily proven using inequality (7). In addition, $S(p\|q)=0$ if and only if $p_{n}=q_{n}$ for all $n$. Thus, it is tempting to interpret $S(p\|q)$ as a metric. However, the relative entropy is neither symmetric, $S(p\|q)\neq S(q\|p)$, nor does it satisfy the triangle inequality, therefore, it should be understood as a measure of distinguishability of the probability distribution $p_{n}$ from a reference distribution $q_{n}$, rather than a distance between them.
@@ VIII Entanglement entropy in DIS
The solution to this equation, with the initial condition $p_{n}(0)=\delta_{n1}$ for $n=1,2,\ldots$ is given by the distribution:
@@
which is normalized to one. The mean multiplicity rises exponentially with rapidity,
$\langle n\rangle={\rm{e}}^{\alpha y}$ (63)
@@
With this result, entropy (58) reads:
$S(y)=\langle n\rangle\ln\langle n\rangle-\left(\langle n\rangle-1\right)\ln\left(\langle n\rangle-1\right)$ (65)
@@
$S_{B}=S_{\text{h}adron}=-\sum_{N}P(N)\ln(N)$ (71)
where $N$ is the number of hadrons (mostly charged pions) in the final state. Hence, one can identify the partonic and hadronic entropy
@@ IX Conclusions
We presented the concept of partonic entropy of the proton, defined in analogy to the Gibbs-Shanon entropy formula for a probability distribution, using the collinear parton distribution functions (PDFs). In this formulation, the probability distribution $P(x)$ is given by the sum of the quark-singlet and gluon distributions multiplied by the Bjorken variable $x$, which is a positive quantity. With such a definition, the momentum sum rule (3) is interpreted as the normalization condition for the probability distribution $P(x)$, which allows to define the non-negative partonic entropy (5).

### v3

@@ II Partonic entropy
Following the Gibbs-Shannon formula [32, 33] for entropy of a probability distribution $\{P_{n}\}$, defined on a discrete set of states nn [^1] In quantum mechanics, $S$ is the von Neumann entropy, $S=-{\rm Tr\,}\rho\ln\rho$, computed in the basis that diagonalizes the density operator $\rho$. ,
@@ V Relative entropy
which can be easily proven using inequality (7):
$\sum_{n}p_{n}\ln\!\left(\frac{p_{n}}{q_{n}}\right)\geq\sum_{n}(p_{n}-q_{n})=0$ (35)
where we used the normalization of the probability distributions to unity. In addition, $S(p\|q)=0$ if and only if $p_{n}=q_{n}$ for all $n$. Thus, it is tempting to interpret $S(p\|q)$ as a metric. However, the relative entropy is neither symmetric, $S(p\|q)\neq S(q\|p)$, nor does it satisfy the triangle inequality, therefore, it should be understood as a measure of distinguishability of the probability distribution $p_{n}$ from a reference distribution $q_{n}$, rather than a distance between them.
@@ VIII Entanglement entropy in DIS
The solution to this equation, with the initial condition $p_{n}(0)=\delta_{n1}$ is given by the distribution:
@@
where n=1,2,…, which is normalized to one,
$\sum_{n=1}^{\infty}p_{n}(y)=1$ (64)
The mean multiplicity rises exponentially with rapidity,
$\langle n\rangle=\sum_{n=1}^{\infty}np_{n}(y)={\rm{e}}^{\alpha y}$ (65)
@@
With this result, the entropy reads:
$S(y)$ $=-\sum_{n=1}^{\infty}p_{n}(y)\ln p_{n}(y)$
$=\langle n\rangle\ln\langle n\rangle-\left(\langle n\rangle-1\right)\ln\left(\langle n\rangle-1\right)$ (67)
@@
$S_{B}=S_{\text{h}adron}=-\sum_{N}P_{N}(y)\ln P_{N}(y)$ (73)
where $P_{N}(y)$ is a probability distribution of $N$ hadrons (mostly charged pions) in the final state at rapidity $y$. Hence, one can identify the partonic and hadronic entropy
@@ IX Conclusions
We presented the concept of partonic entropy of the proton, defined in analogy to the Gibbs-Shannon entropy formula for a probability distribution, using the collinear parton distribution functions (PDFs). In this formulation, the probability distribution $P(x)$ is given by the sum of the quark-singlet and gluon distributions multiplied by the Bjorken variable $x$, which is a positive quantity. With such a definition, the momentum sum rule (3) is interpreted as the normalization condition for the probability distribution $P(x)$, which allows to define the non-negative partonic entropy (5).